OpenStudy (anonymous):

Two companies working together can clear a parcel of land in 20 hours. Working alone, it would take Company A 2 hours longer to clear the land than it would Company B. How long would it take Company B to clear the parcel of land alone? (Round your answer to the nearest tenth.)

OpenStudy (misty1212):

you really do have a lot don't you?

OpenStudy (anonymous):

I have 5 left i think. I answered 3 correctly before this one! :) and for this one, I got a decimal so I don't think it is right!

OpenStudy (misty1212):

i get something bad too but maybe i did it wrong i solved \[\frac{1}{x}+\frac{1}{x+2}=\frac{1}{20}\]

OpenStudy (misty1212):

i got for the slower one \(19+\sqrt{401}\)

OpenStudy (anonymous):

Right and when you multiply both sides by the LCM I got 40x+40=x^2+2x

OpenStudy (anonymous):

Then x^2-38x-40=0

OpenStudy (misty1212):

ok if you used the same equation, then is is probably right, not a pretty answer

OpenStudy (anonymous):

Okay I will try it!